When 12.0 mL of a 3.41×10-4 M sodium hydroxide solution is combined with 12.0 mL of a 4.85×10-4 M aluminum nitrate solution does a precipitate form? ___ (yes or no)

For these conditions the Reaction Quotient, Q, is equal to ___ .

(The given Ksp for Al(OH)3 is 1.9e-33)
Answer #1

Calculate the final volume = 12 +12 = 24 mL Moles of OH = 3.41x10-4 x12x10-3 = 4.092 x 10-6 moles -= 1.705 x 10-4 M Concentra